### v1

@@ 1 Introduction
The proposed paradigm revolutionizes the workflow of pretraining for computational pathology, which allows the pathology FM to possess powerful whole-slide abilities. To the best of our knowledge, this is the first endeavor to inject multimodal knowledge at the slide-level context into pathology foundation models, broadening the scope of contextual modelling from unimodal to multimodal knowledge and from patch-level to slide-level. To systematically investigate the capabilities of mSTAR including unimodal and multimodal applications, we evaluate 7 types of applications on 43 diverse subtasks, resulting in the largest spectrum of downstream tasks. The experimental results have demonstrated that, mSTAR showcased consistent superiority in not only unimodal but also multimodal applications, with statistically significant differences compared to previous state-of-the-art (SOTA) FMs (Figure 1f-g). These findings indicate the noteworthy power of multimodal knowledge at the whole-slide context, which endows pathology foundation models with enhanced capabilities that extend beyond the patch-level unimodal context, enabling them to achieve more generalized performance.
@@ 2.2 Pathological Slide Classification for Diagnosis and Treatment
When considering the incorporation with the pretrained aggregator, mSTAR+ exhibited further advancements on various datasets and achieved the best on 5 out of 6 diagnostic tasks. For example, compared to the second-best compared FM, mSTAR+ outperformed by +2.01% on CAMELYON ($P<0.001$). Furthermore, an interesting finding from Figure 3e (Extended Data Table 4) was that the pretrained aggregator was even able to collaborate with other patch extractors achieving critical performance gains, although it was solely pretrained on features extracted by mSTAR. As shown in Figure 3e, it consistently displayed superior performance in the majority of tasks for different FMs. Among these, the pretrained aggregator achieved a notable increase of up to +14.94% ($P<0.001$) compared to the one built from scratch, although there were slight decreases (-1.15% at most) of performance in some specific subtasks. From Figure 3e, we also observed that the enhancements made to CONCH, UNI, and mSTAR were not substantial, which could be attributed to the fact that their performances have already approached 100% and the room for improvement was significantly limited on these 6 subtasks.
@@
From the perspective of consistency, mSTAR with ABMIL achieved the best performance on 5 out of 6 tasks (Figure 3c and Extended Data Table 2), in which 4 tasks exhibited a significance difference ($P<0.001$). With TransMIL, clearer improvements (Figure 3d and Extended Data Table 3) were observed that 4 out of 6 tasks demonstrated a significant difference ($P<0.05$). Taking a closer look at Macro-AUC, mSTAR outperformed other SOTA foundation models by +2.29% overall with a critical difference ($P<0.001$) on average (Extended Data Table 2-3). Specifically, for TransMIL, a substantial increase of +3.71% on average Macro-AUC across 6 molecular tasks was observed, although mSTAR achieved slight improvement with ABMIL ($P<0.05$). In particular, substantial performance increases of +4.60% and +4.21% were observed on BRCA-Molecular and CRC-Molecular tasks ($P<0.001$), respectively.
@@ 2.3 Pathological Survival Analysis for Prognosis
Figure 4: Pathological Survival Analysis Performance. a, Comparison of C-Index between mSTAR and compared methods on 9 TCGA datasets. They are evaluated with ABMIL and TransMIL trained from scratch, based on different patch extractors. “Overall” refers to the average results between ABMIL and TransMIL. b, The average C-Index on 9 TCGA datasets. c-d The survival analysis performance (C-Index and 95% CI) on 9 TCGA datasets with ABMIL and TransMIL, respectively. If the performance of mSTAR is the best one compared against the compared FMs, P-value would be presented. mSTAR+ represents mSTAR (patch extractor) equipped with the pretrained aggregator (TransMIL+). The colors of legends are shared across sub-figures a-d. e The performance (C-Index) gains of the aggregator (TransMIL) between ’before’ and ’after’ pretraining among 9 TCGA cancer types. ’Positive’ values indicate performance increases compared with the ones equipped with TransMIL built from scratch, and vice versa. All of performance increases showed significant differences with $P<0.05$, except the ones marked by $\times$.
Prognostic analysis is an intricate clinical endeavor, necessitating a thorough analysis from a multitude of facets. In this regard, the integration of multimodal data has proven instrumental in enabling more comprehensive prognostic assessments [8, 9, 10]. Therefore, it is crucial to delve into the exploration of the role of multimodal knowledge within the broader whole-slide context in enhancing prognosis estimation. In this study, we assessed a key prognostic task, overall survival (OS) prediction, on top of pathological tissue slides.
mSTAR advances Cancer Survival Prediction. We investigate the capability of predicting overall survival (OS) across 9 diverse cancers, based on the metric of C-Index, a metric commonly used in survival prediction. To avoid data leakage, the data used for evaluation were held out from pretraining data. To ensure the reliable evaluation, we excluded cancer types with fewer than 450 cases or those with too few uncensored data out of the 32 cancer datasets in TCGA [19], resulting in 9 cancer datasets: Breast Invasive Carcinoma (BRCA), Colon Adenocarcinoma and Rectum Adenocarcinoma (CRC), Glioblastoma Multiforme and Brain Lower Grade Glioma (GBMLGG), Head and Neck Squamous Cell Carcinoma (HNSC), Kidney Renal Clear Cell Carcinoma (KIRC), Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Skin Cutaneous Melanoma (SKCM) and Uterine Corpus Endometrial Carcinoma (UCEC). Similarly, following the settings of slide-level classification, we also used ABMIL and TransMIL for evaluating patch extractors and then the performance of the pretrained slide aggregator was assessed. More details are presented in Section 4.3.
@@
When specifically assessing the C-Index metric, the average C-Index showcased significant performance gains compared against the second-best model (UNI) by +1.98% for TransMIL ($P<0.005$), +0.53% for ABMIL ($P<0.05$), and +1.30% for the average performance of different MIL backbones (Overall, $P<0.001$). Next, we delved into details of performances across 9 datasets, where we observed apparent improvements ($P<0.001$) of +3.87% on UCEC, +3.34% on LUAD, +2.89% on LUSC and +1.46% on SKCM. These results indicated the effectiveness of multimodal knowledge involved in mSTAR in extracting discriminative features for distinguishing complex survival patterns in pathological images.
We also investigated the performance of incorporating the pretrained aggregator. From observations of Figure 4d, mSTAR+ equipped with the pretrained aggregator pushed performance to higher levels (+5.23% on SKCM, +4.9% on UCEC, +3.0% on BRCA, +2.69% on KIRC and +2.33% on CRC, $P<0.001$) compared to the second-performing approaches. By comparing TransMIL “before” and “after” pretraining for various extractors (Figure 4e, Extended Data Table 7), we observed that, in the majority of cases, the pretrained aggregator exhibited superior performance, surpassing the one built from scratch by a margin of up to +7.79% ($P<0.001$). This indicates the introduction of pathology report and gene expression profiles is beneficial to survival analysis by providing additional expert knowledge and molecular information. In addition, a similar trend to that of slide classification was observed that, PLIP and CONCH that had a lack of knowledge of gene expression profiles exhibited more notable improvements than mSTAR, when collaborating with our aggregator pretrained on gene data. This supports the analysis illustrated in the results of slide classification again, in which multimodal knowledge has been transferred into the patch extractor mSTAR, leading to relatively moderate improvements when using the pretrained aggregator. In other words, mSTAR itself has absorbed the knowledge in pathology reports and gene expression profiles, which has contributed to survival prediction.
@@ 2.4 Multimodal Capability
On the whole, mSTAR has clearly outperformed other SOTA methods by a wide margin. Considering average rank, mSTAR ranked between 1.22 and 1.67 among various fusion models and the overall rank is 1.47, which left the second-best approach UNI far behind (Figure 5a) ranking at 2.56 to 3.00 and 2.68 on average. For average C-Index (Figure 5b), mSTAR achieved consistent and notable enhancement in multimodal fusion with a significant difference, with average performance increases of +1.5% ($P<0.001$). Among different multimodal fusion models, mSTAR outperformed the second-best UNI by +1.7% ($P=0.02$) for MCAT, +1.1% ($P=0.01$) for Porpoise, +2.5% ($P<0.005$) for MOTCat and +1.9% ($P<0.005$) for CMTA.
Across various datasets, MCAT surpassed other SOTA approaches on 7 out of 9 tasks (Figure 5c and Extended Data Table 8), especially on LUAD (+3.3%, $P<0.001$) and BRCA (+1.9%, $P<0.001$). In the case of MOTCat, mSTAR excelled in 7 out of 9 tasks (Figure 5d and Extended Data Table 10) with performance increases of 0.6%-4.8% ($P<0.001$). mSTAR with Porpoise demonstrated superior performance in the majority of tasks, topping 5 out of 9 datasets (Figure 5e and Extended Data Table 9), which increased the second-best model by 0.3%-3.8% ($P<0.001$). For CMTA, mSTAR delivered the highest performance in 7 of 9 (Figure 5f and Extended Data Table 11), advancing the second-best one by 0.4%-4.1% ($P<0.001$).
@@
Across 6 subtasks, mSTAR ranked the 1st place overall and achieved the best performance on 4 subtasks (Figure 6b and Extended Data Table 12), where mSTAR outperformed second-best FMs on average by a large margin (+4.1% on Camelyon, +1.1% on UBC-OCEAN, +1.7% on BCNB-PR and +0.9% for overall), as shown in Figure 6c (Extended Data Table 12). To demonstrate the benefits of multimodal knowledge, we also presented the performance increase compared to the vision-only SOTA FM, UNI. On 4 out of 6 subtasks, mSTAR exhibited better performance on average, with an additional increase on PANDA by +1.7%.
@@ Appendix A Extended Data
Diagnosis CAMELYON 0.9236 (0.8905,0.9567) 0.9242 (0.8915,0.9569) 0.9823 (0.9654,0.9992) 0.9814 (0.9683,0.9945) 0.9934 (0.9865,1.000)
PANDA 0.8872 (0.8811,0.8933) 0.9045 (0.8988,0.9102) 0.922 (0.9169,0.9271) 0.9455 (0.9410,0.9500) 0.9468 (0.9423,0.9513)
UBC-OCEAN 0.894 (0.8644,0.9236) 0.9453 (0.9230,0.9676) 0.9724 (0.9606,0.9842) 0.9732 (0.9612,0.9852) 0.9764 (0.9637,0.9891)
@@
Table 4: Macro-AUC of Slide Classification with the pretrained TransMIL on 12 datasets. The performance gains of the aggregator (TransMIL) between ’before’ and ’after’ pretraining. ’Positive’(+) values indicate performance increases ($\uparrow$) compared with the ones equipped with TransMIL built from scratch, and vice versa. The best improvement is bolded.
@@
Table 7: C-Index of Survival Analysis with the pretrained TransMIL on 9 TCGA datasets. The performance gains of the aggregator (TransMIL) between ’before’ and ’after’ pretraining. ’Positive’(+) values indicate performance increases ($\uparrow$) compared with the ones equipped with TransMIL built from scratch, and vice versa. The best improvement is bolded.
@@
Table 12: Average Macro-AUC of Few-shot Slide Classification across different shots on 6 datasets. Best performing model for each metric is in bold. Mean+std is presented.
@@
CAMELYON 0.525±0.0185 0.5457±0.0333 0.5039±0.0202 0.666±0.087 0.7074±0.0634
PANDA 0.5741±0.0203 0.6641±0.0451 0.6484±0.0305 0.6044±0.0219 0.621±0.0295
UBC-OCEAN 0.5893±0.0224 0.7232±0.0118 0.7147±0.0269 0.7635±0.0232 0.7742±0.0317
BCNB-ER 0.5625±0.0251 0.5823±0.0343 0.6027±0.0679 0.6232±0.0485 0.6235±0.0648
BCNB-PR 0.535±0.0493 0.5985±0.0274 0.614±0.0165 0.643±0.0391 0.6602±0.0303
BCNB-HER2 0.5592±0.054 0.5138±0.0139 0.5235±0.034 0.574±0.0386 0.5423±0.0156
@@
Table 13: Macro-AUC of Few-shot Slide Classification on CAMELYON across different shots. Best performing model for each metric is bolded. Mean+std is presented.
@@
Table 14: Macro-AUC of Few-shot Slide Classification on PANDA across different shots. Best performing model for each metric is bolded. Mean+std is presented.
@@
Table 15: Macro-AUC of Few-shot Slide Classification on UBC-OCEAN across different shots. Best performing model for each metric is bolded. Mean+std is presented.
@@
Table 16: Macro-AUC of Few-shot Slide Classification on BCNB-ER across different shots. Best performing model for each metric is bolded. Mean+std is presented.
@@
Table 17: Macro-AUC of Few-shot Slide Classification on BCNB-PR across different shots. Best performing model for each metric is bolded. Mean+std is presented.
@@
Table 18: Macro-AUC of Few-shot Slide Classification on BCNB-HER2 across different shots. Best performing model for each metric is bolded. Mean+std is presented.
@@
CAMELYON 0.541 (0.4548,0.6272) 0.622 (0.5338,0.7102) 0.638 (0.5498,0.7262) 0.685 (0.6007,0.7693) 0.742 (0.6675,0.8165)
PANDA 0.48 (0.4663,0.4937) 0.561 (0.5473,0.5747) 0.582 (0.5663,0.5977) 0.45 (0.4343,0.4657) 0.462 (0.4463,0.4777)
UBC-OCEAN 0.445 (0.3686,0.5214) 0.762 (0.6914,0.8326) 0.37 (0.3073,0.4327) 0.694 (0.6195,0.7685) 0.692 (0.6214,0.7626)
BCNB-ER 0.631 (0.535,0.727) 0.563 (0.4728,0.6532) 0.5 (0.5,0.5) 0.618 (0.5239,0.7121) 0.725 (0.6407,0.8093)
BCNB-PR 0.554 (0.458,0.65) 0.523 (0.4348,0.6112) 0.5 (0.5,0.5) 0.682 (0.5977,0.7663) 0.722 (0.6436,0.8004)
BCNB-HER2 0.549 (0.4549,0.6431) 0.516 (0.4298,0.6022) 0.518 (0.4278,0.6082) 0.465 (0.3729,0.5571) 0.435 (0.3448,0.5252)
@@
Table 21: Data splits (cases) of 9 cancer datasets on TCGA. COADREAD=COAD+READ and GBMLGG=GBM+LGG. Apart from these specific cancer types, other TCGA datasets were used for pretraining without data partitioning.

### v2

@@ 1 Introduction
The proposed paradigm revolutionizes the workflow of pretraining for computational pathology, which allows the pathology FM to possess powerful whole-slide abilities. Furthermore, to the best of our knowledge, this is the first endeavor to inject multimodal knowledge at the slide-level context into pathology foundation models, broadening the scope of contextual modelling from unimodal to multimodal knowledge and from patch-level to slide-level. To systematically investigate the capabilities of mSTAR including unimodal and multimodal applications, we evaluate 7 types of applications on 43 diverse subtasks, resulting in the largest spectrum of downstream tasks. The experimental results have demonstrated that, mSTAR showcased consistent superiority in not only unimodal but also multimodal applications on average, with statistically significant differences compared to previous state-of-the-art (SOTA) FMs (Figure 1f-g). These findings indicate the noteworthy power of multimodal knowledge at the whole-slide context, which endows pathology foundation models with enhanced capabilities that extend beyond the patch-level unimodal context, enabling them to achieve more generalized performance.
@@ 2.2 Pathological Slide Classification for Diagnosis and Treatment
When considering the incorporation with the pretrained aggregator, mSTAR+ exhibited further advancements on various datasets and achieved the best on 5 out of 6 diagnostic tasks. For example, compared to the second-best compared FM, mSTAR+ outperformed by +2.01% on CAMELYON ($P<0.001$). Furthermore, an interesting finding from Figure 3e (Extended Data Table 4) was that the pretrained aggregator was even able to collaborate with other patch extractors achieving critical performance gains, although it was solely pretrained on features extracted by mSTAR. As shown in Figure 3e, it consistently displayed superior performance in the majority of tasks for different FMs. Among these, the pretrained aggregator achieved a notable increase of up to +9.10% ($P<0.001$) compared to the one built from scratch, although there were slight decreases (-1.15% at most) of performance in some specific subtasks. From Figure 3e, we also observed that the enhancements made to CONCH, UNI, and mSTAR were not substantial, which could be attributed to the fact that their performances have already approached 100% and the room for improvement was significantly limited on these 6 subtasks.
@@
From the perspective of consistency, mSTAR with ABMIL achieved the best performance on 5 out of 6 tasks (Figure 3c and Extended Data Table 2), in which 4 tasks exhibited a significance difference ($P<0.001$). With TransMIL, clearer improvements (Figure 3d and Extended Data Table 3) were observed that 4 out of 6 tasks demonstrated a significant difference ($P<0.05$). Taking a closer look at Macro-AUC, mSTAR outperformed other SOTA foundation models by +2.23% overall with a critical difference ($P<0.001$) on average across different MIL backbones (Extended Data Table 2-3). Specifically, for TransMIL, a substantial increase of +3.71% on average Macro-AUC across 6 molecular tasks was observed, although mSTAR achieved slight improvement with ABMIL ($P<0.05$). In particular, substantial performance increases of +4.64% and +4.21% were observed on BRCA-Molecular and CRC-Molecular tasks ($P<0.001$), respectively.
@@ 2.3 Pathological Survival Analysis for Prognosis
Figure 4: Pathological Survival Analysis Performance. a, Comparison of C-Index between mSTAR and compared methods on 9 TCGA datasets. They are evaluated with ABMIL and TransMIL trained from scratch, based on different patch extractors. “Overall” refers to the average results between ABMIL and TransMIL. b, The average C-Index on 9 TCGA datasets. c-d The survival analysis performance (C-Index and 95% CI) on 9 TCGA datasets with ABMIL and TransMIL, respectively. If the performance of mSTAR is the best one compared against the compared FMs, P-value would be presented. mSTAR+ represents mSTAR (patch extractor) equipped with the pretrained aggregator (TransMIL+). The colors of legends are shared across sub-figures a-d. e The performance (C-Index) gains of the aggregator (TransMIL) between “before” and “after” pretraining among 9 TCGA cancer types. “Positive” values indicate performance increases compared with the ones equipped with TransMIL built from scratch, and vice versa. All of performance increases showed significant differences with $P<0.05$, except the ones marked by $\times$.
Prognostic analysis is an intricate clinical endeavor, necessitating a thorough analysis from a multitude of facets. In this regard, the integration of multimodal data has proven instrumental in enabling more comprehensive prognostic assessments [8, 9, 10, 23]. Therefore, it is crucial to delve into the exploration of the role of multimodal knowledge within the broader whole-slide context in enhancing prognosis estimation. In this study, we assessed a key prognostic task, overall survival (OS) prediction, on top of pathological tissue slides.
mSTAR advances Cancer Survival Prediction. We investigate the capability of predicting overall survival (OS) across 9 diverse cancers, based on the metric of C-Index, a metric commonly used in survival prediction. To avoid data leakage, the data used for evaluation were held out from pretraining data. To ensure the reliable evaluation, we excluded cancer types with fewer than 400 cases or those with too few uncensored data out of the 32 cancer datasets in TCGA [19], resulting in 9 cancer datasets: Breast Invasive Carcinoma (BRCA), Colon Adenocarcinoma and Rectum Adenocarcinoma (CRC), Glioblastoma Multiforme and Brain Lower Grade Glioma (GBMLGG), Head and Neck Squamous Cell Carcinoma (HNSC), Kidney Renal Clear Cell Carcinoma (KIRC), Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Skin Cutaneous Melanoma (SKCM) and Uterine Corpus Endometrial Carcinoma (UCEC). Similarly, following the settings of slide-level classification, we also used ABMIL and TransMIL for evaluating patch extractors and then the performance of the pretrained slide aggregator was assessed. More details are presented in Section 4.3.
@@
When specifically assessing the C-Index metric, the average C-Index showcased significant performance gains compared against the second-best model (UNI) by +1.98% for TransMIL ($P<0.005$), +0.53% for ABMIL ($P<0.05$), and +1.30% for the average performance of different MIL backbones (Overall, $P<0.001$). Next, we delved into details of performances across 9 datasets, where we observed apparent improvements ($P<0.001$) of +3.7% on SKCM, +3.34% on LUAD, +2.87% on LUSC and +1.03% on UCEC over the second-best model based on TransMIL (4d and Extended Data Table 6). These results indicated the effectiveness of multimodal knowledge involved in mSTAR in extracting discriminative features for distinguishing complex survival patterns in pathological images.
We also investigated the performance of incorporating the pretrained aggregator. From observations of Figure 4d, mSTAR+ equipped with the pretrained aggregator pushed performance to higher levels (+5.23% on SKCM, +4.9% on UCEC, +2.96% on BRCA, +2.60% on KIRC and +2.33% on CRC, $P<0.001$) compared to the second-performing approaches. By comparing TransMIL “before” and “after” pretraining for various extractors (Figure 4e, Extended Data Table 7), we observed that, in the majority of cases, the pretrained aggregator exhibited superior performance, surpassing the one built from scratch by a margin of up to +7.79% ($P<0.001$). This indicates the introduction of pathology report and gene expression profiles is beneficial to survival analysis by providing additional expert knowledge and molecular information. In addition, a similar trend to that of slide classification was observed that, PLIP and CONCH that had a lack of knowledge of gene expression profiles exhibited more notable improvements than mSTAR, when collaborating with our aggregator pretrained on gene data. This supports the analysis illustrated in the results of slide classification again, in which multimodal knowledge has been transferred into the patch extractor mSTAR, leading to relatively moderate improvements when using the pretrained aggregator. In other words, mSTAR itself has absorbed the knowledge in pathology reports and gene expression profiles, which has contributed to survival prediction.
@@ 2.4 Multimodal Capability
On the whole, mSTAR has clearly outperformed other SOTA methods by a wide margin. Considering average rank, mSTAR ranked between 1.22 and 1.67 among various fusion models and the overall rank is 1.47, which left the second-best approach UNI far behind (Figure 5a) ranking at 2.56 to 3.00 and 2.68 on average. For average C-Index (Figure 5b), mSTAR achieved consistent and notable enhancement in multimodal fusion with a significant difference, with average performance increases of +1.8% ($P<0.001$). Among different multimodal fusion models, mSTAR outperformed the second-best UNI by +1.7% ($P=0.02$) for MCAT, +1.1% ($P=0.01$) for Porpoise, +2.5% ($P<0.005$) for MOTCat and +1.9% ($P<0.005$) for CMTA.
Across various datasets, based on MCAT, mSTAR surpassed other SOTA approaches on 7 out of 9 tasks (Figure 5c and Extended Data Table 8), especially on LUAD (+3.3%, $P<0.001$) and BRCA (+1.9%, $P<0.001$). In the case of MOTCat, mSTAR excelled in 7 out of 9 tasks (Figure 5d and Extended Data Table 10) with performance increases of 0.6%-4.8% ($P<0.001$). mSTAR with Porpoise demonstrated superior performance in the majority of tasks, topping 5 out of 9 datasets (Figure 5e and Extended Data Table 9), which increased the second-best model by up to +3.8% ($P<0.001$). For CMTA, mSTAR delivered the highest performance in 7 of 9 (Figure 5f and Extended Data Table 11), advancing the second-best one by 0.4%-4.1% ($P<0.001$).
@@
Across 6 subtasks, mSTAR ranked the 1st place overall and achieved the best performance on 4 subtasks (Figure 6b and Extended Data Table 12), where mSTAR outperformed second-best FMs on average by a large margin (+4.1% on Camelyon, +1.0% on UBC-OCEAN, +2.0% on BCNB-PR and +0.9% for overall), as shown in Figure 6c (Extended Data Table 12). To demonstrate the benefits of multimodal knowledge, we also presented the performance increase compared to the vision-only SOTA FM, UNI. On 4 out of 6 subtasks, mSTAR exhibited better performance on average, with an additional increase on PANDA by +1.7%.
@@ Appendix A Extended Data
Diagnosis CAMELYON 0.9236 (0.8905,0.9567) 0.9242 (0.8915,0.9569) 0.9823 (0.9654,0.9992) 0.9814 (0.9683,0.9945) 0.9934 (0.9865,1.0000)
PANDA 0.8872 (0.8811,0.8933) 0.9045 (0.8988,0.9102) 0.9220 (0.9169,0.9271) 0.9455 (0.9410,0.9500) 0.9468 (0.9423,0.9513)
UBC-OCEAN 0.8940 (0.8644,0.9236) 0.9453 (0.9230,0.9676) 0.9724 (0.9606,0.9842) 0.9732 (0.9612,0.9852) 0.9764 (0.9637,0.9891)
@@
Table 4: Macro-AUC of Slide Classification with the pretrained TransMIL on 12 datasets. The performance gains of the aggregator (TransMIL) between “before” and “after” pretraining. “Positive”(+) values indicate performance increases ($\uparrow$) compared with the ones equipped with TransMIL built from scratch, and vice versa. The best improvement is bolded.
@@
Table 7: C-Index of Survival Analysis with the pretrained TransMIL on 9 TCGA datasets. The performance gains of the aggregator (TransMIL) between “before”” and “after” pretraining. ’Positive’(+) values indicate performance increases ($\uparrow$) compared with the ones equipped with TransMIL built from scratch, and vice versa. The best improvement is bolded.
@@
Table 12: Average Macro-AUC of Few-shot Slide Classification across different shots on 6 datasets. Best performing model for each metric is in bold. Mean±std is presented.
@@
CAMELYON 0.525±0.019 0.546±0.033 0.504±0.020 0.666±0.087 0.707±0.063
PANDA 0.574±0.020 0.664±0.045 0.648±0.031 0.604±0.022 0.621±0.030
UBC-OCEAN 0.589±0.022 0.723±0.012 0.715±0.027 0.764±0.023 0.774±0.032
BCNB-ER 0.563±0.025 0.582±0.034 0.603±0.068 0.623±0.049 0.624±0.065
BCNB-PR 0.535±0.049 0.599±0.027 0.614±0.017 0.640±0.039 0.660±0.030
BCNB-HER2 0.559±0.054 0.514±0.014 0.524±0.034 0.574±0.039 0.542±0.016
@@
Table 13: Macro-AUC of Few-shot Slide Classification on CAMELYON across different shots. Best performing model for each metric is bolded. Mean±std is presented.
@@
Table 14: Macro-AUC of Few-shot Slide Classification on PANDA across different shots. Best performing model for each metric is bolded. Mean±std is presented.
@@
Table 15: Macro-AUC of Few-shot Slide Classification on UBC-OCEAN across different shots. Best performing model for each metric is bolded. Mean±std is presented.
@@
Table 16: Macro-AUC of Few-shot Slide Classification on BCNB-ER across different shots. Best performing model for each metric is bolded. Mean±std is presented.
@@
Table 17: Macro-AUC of Few-shot Slide Classification on BCNB-PR across different shots. Best performing model for each metric is bolded. Mean±std is presented.
@@
Table 18: Macro-AUC of Few-shot Slide Classification on BCNB-HER2 across different shots. Best performing model for each metric is bolded. Mean±std is presented.
@@
CAMELYON 0.541 (0.455,0.627) 0.622 (0.534,0.710) 0.638 (0.550,0.726) 0.685 (0.601,0.769) 0.742 (0.668,0.817)
PANDA 0.480 (0.466,0.494) 0.561 (0.547,0.575) 0.582 (0.566,0.598) 0.450 (0.434,0.466) 0.462 (0.446,0.478)
UBC-OCEAN 0.445 (0.369,0.521) 0.762 (0.691,0.833) 0.370 (0.307,0.433) 0.694 (0.620,0.769) 0.692 (0.621,0.763)
BCNB-ER 0.631 (0.535,0.727) 0.563 (0.473,0.653) 0.500 (0.500,0.500) 0.618 (0.524,0.712) 0.725 (0.641,0.809)
BCNB-PR 0.554 (0.458,0.650) 0.523 (0.435,0.611) 0.500 (0.500,0.500) 0.682 (0.598,0.766) 0.722 (0.644,0.800)
BCNB-HER2 0.549 (0.455,0.643) 0.516 (0.430,0.602) 0.518 (0.428,0.608) 0.465 (0.373,0.557) 0.435 (0.345,0.525)
@@
Table 21: Data splits (cases) of 9 cancer datasets on TCGA. CRC=COAD+READ and GBMLGG=GBM+LGG. Apart from these specific cancer types, other TCGA datasets were used for pretraining without data partitioning.